OpenStudy (anonymous):

HELP HELP HELP HELP , Tracy Needs To Find The Length And Width Of The Equipment Storage Room Where She Works. The Room Is 1 Meter Longer Than It Is Wide. Its Perimeter Is 14 Meters. Find The Width In Meters.

OpenStudy (anonymous):

 L + L + W + W = 14

OpenStudy (anonymous):

L + L + W + W = 14 L = W + 1 W + 1 + W + 1 + W + W = 14 can you do the rest?

OpenStudy (anonymous):

No I Really Need The Help

OpenStudy (anonymous):

The perimeter is 14 so L + L + W + W = 14 You get that?

OpenStudy (anonymous):

So Its 14

OpenStudy (anonymous):

It's 1 meter longer than it's wide so L = W + 1 Get that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So, now i write W+1 instead of L in the first equation: ( W + 1 ) + ( W + 1 ) + W + W = 14 get that?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Now we add it all up and get the answer W + W + W + W + 1 + 1 = 14 4W + 2 = 14 4W = 14 - 2 4W = 12 W = 12/4 W = 3 And if W = 3 then L = 3+1 = 4